# Ultrasound matrix imaging for transcranial in-vivo localization microscopy

### v1

@@
Ultrasound matrix imaging for transcranial in-vivo localization microscopy
Transcranial ultrasound imaging is usually limited by skull-induced attenuation and high-order aberrations. By using contrast agents such as microbubbles in combination with ultrafast imaging, not only can the signal-to-noise ratio be improved, but super-resolution images down to the micrometer scale of the brain vessels can be obtained. However, ultrasound localization microscopy (ULM) remains impacted by wave-front distortions that limit the microbubble detection rate and hamper their localization. In this work, we show how matrix imaging, which relies on the prior recording of the reflection matrix, can provide a solution to those fundamental issues. As an experimental proof-of-concept, an in-vivo reconstruction of deep brain microvessels is performed on three anesthetized sheeps. The compensation of wave distortions is shown to drastically enhance the contrast and resolution of ULM. This experimental study thus opens up promising perspectives for a transcranial and non-ionizing observation of human cerebral microvascular pathologies, such as stroke.
@@ INTRODUCTION
To solve the fundamental problem of aberrations, adaptive focusing was introduced years ago in astronomy to compensate for phase distortions undergone by starlight when it passes through the atmosphere (?, ?, ?, ?). This idea was then transposed to ultrasound imaging by tailoring a focusing law that either optimizes an image metrics (?, ?) or maximizes the spatial correlations of back-scattered echoes (?, ?, ?, ?, ?). In the specific case of ULM, several methods have been proposed to correct the aberrations caused by the skull. A first strategy (?) is based on deep learning algorithms trained with numerical simulations. Although it gives impressive in-vivo results for small animals, it is hampered by the specificity of the training set, so its transfer to other experimental configurations (probe, organ, SNR) may require tedious training procedures. The second method (?) uses microbubbles as ultrasound guide stars but suffers several drawbacks. On the one hand, microbubbles should be detected and isolated enough, which can be extremely difficult to assess because of the strongly distorted focal spot across the skull, On the other hand, the number of detected micro-bubbles should be large enough for self-averaging in each isoplanatic patch.
Interestingly, inspired by previous seminal works (?, ?, ?), a matrix approach of ultrasound imaging has been recently developed to overcome aberrations without any guide star (?, ?). Based on the recording of the reflection associated with the ultrasonic probe and the linearity of the wave equation, it allows the simulation of wave focusing in post-processing and an iteration of this process towards local adaptive focusing laws in ultrasonic speckle (?). An optimized contrast and close-to-ideal resolution is then obtained for each voxel of the ultrasound image. UMI has been successfully applied to transcranial 3D imaging of a head phantom (?).
The goal of this paper is to demonstrate the interest of UMI for in-vivo brain imaging. In particular, we will show how: (i) UMI can be used to quantify aberrations and multiple scattering in in-vivo transcranial imaging; (ii) UMI can exploit static speckle from brain tissue to tailor complex focusing laws; (iii) UMI can leverage these focusing laws to increase the detection rate of microbubbles in ULM and improve their localization; (iv) UMI can be fruitfully combined to ULM in order to provide a three-dimensional, artifacts-free, super-resolved and contrasted image of brain vessels.
@@ RESULTS
In this paper, we demonstrate transcranial brain imaging by overcoming the major challenges of: (i) skull bone aberration with 3D UMI; (ii) diffraction-limited resolution with the super-resolution feature provided by ULM. As a proof-of-concept, the same experimental protocol (Fig. 1) has been applied to three anesthetized sheeps (Methods). A array of transducers is placed on the shaved head of each sheep directly over the crest of the frontal bone (Fig. 1B). The probe is driven in 1-2 frequency bandwidth in order to limit the attenuation by skull bone (?). The first ultrasound sequence consists in recording a high-dimension reflection matrix (Fig. 1C) that contains the scattered wave-field recorded by each transducer of the probe as a function of the echo time for a set of 324 incident diverging waves. This illumination basis is associated with a set of virtual transducers placed behind the probe (see Methods and Fig. S1). It maximizes the signal-to-noise ratio across the targeted field-of-view (?).
@@
where and are the coordinates of the input and output focal points, respectively. and are the expected travel times for the wave to go from the source/detector to the input/output focusing point (Methods, Eq. 4). The result of Eq. 1 is a focused reflection matrix whose coefficients probe the cross-talk between a virtual source and detector at lateral positions and , respectively, both located at the same depth inside the medium (Fig. 2A). This matrix enables a local quantification of the focusing quality and multiple scattering by providing a local point spread function in reflection (RPSF) (?, ?, ?) (Methods, Eq. 7). Figures 3C and F display maps of RPSFs at different depths in sheeps n5 and 6. In the single scattering regime and in absence of aberrations, the RPSF would exhibit a confocal diffraction-limited peak. On the contrary, the various RPSFs here display a distorted focal spot induced by wave velocity heterogeneities on top of an incoherent background due to multiple scattering events induced by the skull. Strikingly, whatever the imaging depth, the RPSF looks more distorted and less contrasted for sheep n6 (Fig. 3F) than for sheep n5 (Fig. 3C).
On the one hand, the aberration level can be quantified by investigating the spatial extent of the RPSF (?, ?) (Figure S2). The depth evolution of is shown in Fig. 3D. A general trend is a degradation of the resolution with depth as predicted by diffraction theory but also a more important aberration level in sheep n6. This difference of behavior can be understood by looking at the micro-CT of each skull (Figs. 3A,B). While the skull of sheep n5 shows a regular thickness mm, skull n6 displays a more irregular shape: mm (?). This difference of morphology between the two skulls can explain the higher degree of aberrations observed in the latter case.
On the other hand, the multiple scattering rate can be estimated from the incoherent background of the RPSF (Methods, Eq. 9). It appears that the multiple scattering rate is far from being negligible even after the beamforming process since it reaches a value of 30 % in the mm-region for sheep n6 ((Fig. 3E). As the aberration level, the multiple scattering rate is larger for this sheep over almost the whole depth range. This feature can be understood by looking at the internal structure of the bone. Indeed, the diploë volume ratio is much higher in sheep n6 (%) than in sheep n5 (%) (?). This high degree of heterogeneity seems to account for the higher multiple scattering rate observed in sheep n6.
@@
The complex speed-of-sound distribution highlighted by Fig. 3B hampers ultrasound brain imaging. The confocal image of the brain, that can be extracted from the diagonal coefficients of (Methods, Eq. 5), is blurred by the skull heterogeneities (Fig. 4A). Fortunately, an optimized contrast and diffraction-limited resolution can be recovered using the distortion matrix concept (?, ?, ?). This matrix contains the wave distortions exhibited by each reflected wave-front in the transducer basis (Fig. 2C). The choice of this basis is dictated by the fact that most aberrations are induced by the skull heterogeneities, i.e in the vicinity of the probe. An iterative phase reversal process is then applied to the distortion matrix to provide an estimation of an aberration phase law, , at each depth (Methods, Fig. 2D and Movie S1). Examples of aberration phase laws are displayed in Fig. 4C. They show a complex feature, as quantified by the Strehl ratio that provides the ratio between the energy at focus with and without aberrations (?) (Methods, Eq. 18). For instance, the aberration law measured at depth mm exhibits a Strehl ratio . This extremely low value illustrates the detrimental impact of skull heterogeneities on the focusing process. Those aberration phase laws can be leveraged to compensate for the wave distortions exhibited by the focused reflection matrix (Methods). The resulting confocal image is displayed in Fig. 4E. It exhibits a clear improvement compared to the initial confocal image (Fig. 4A), with a contrast gain of 8 dB for the deepest bright scatterers ( and mm). The efficiency of the correction process can be assessed by comparing the maps of RPSFs before (Fig. 4B) and after (Fig. 4D) aberration compensation. The spatial resolution is improved by a factor between 2 and 3 and the multiple scattering background is decreased by more than 10 dB beyond mm (Figure S3). Nevertheless, the obtained RPSFs remain imperfect in Fig. 4D. High-order aberrations subsist because of their anisoplanicity. Last but not least, even though the confocal image shows a clear improvement, the diffraction-limited resolution remains limited by the probe size : .
@@
A deconvolution of the dynamic image with a theoretical PSF and the application of a detection threshold provides a localization of bubbles in the field-of-view (Methods). The position of the detected bubbles are highlighted with red disks in Fig. 5A and B. While several microbubbles are wrongly assigned to the high side-lobes of each distorted focal spot in Fig. 5A, UMI drastically reduces this number of false detection in Fig. 5B by restoring a satisfying PSF, in particular in the colored squares. Despite this reduction of false alarms, the number of localized microbubbles is increased. Figure 5C illustrates this assertion by showing, for sheep n, the higher number of detected bubbles over time with UMI. The different peaks corresponds to each injection of microbubbles.
To further reduce the false alarm rate and better appreciate the effect of UMI, microbubbles can be tracked over time (Methods). Fig. 5D shows the increase of detected tracks provided by UMI as a function of their length. UMI significantly increases the number of detected tracks by a factor ranging from 40 to 250% according to the track length. Table 1 shows the same observable for different minimum path lengths. The general trend is the same for each acquisition, thereby showing the robustness of UMI and its benefit in each imaging configuration.
@@
2 frames | >3 frames | >5 frames | >10 frames | >30 frames | ||
@@
However, the skull heterogeneities also induce artifacts such as the duplication of one artery that can be noticed at the bottom of the conventional ULM image (Figs. 6A and C). This type of artifact has already been observed in transcranial imaging in both mouse (?) and human (?). Strikingly, the focusing laws extracted by UMI (Fig. 4C) allow the removal of such artifacts as highlighted by the corrected ULM images in Figs. 6B and D. More generally, the higher detection rate and correct repositioning of micro-bubbles lead to a better contrast for ULM as highlighted by the vessel network on top of the Willis circle (25-35 mm) that appears much brighter (+6 dB) and better resolved. The same resolution and contrast improvement can be observed for the other 3D ULM images shown in Figure S4 and corresponding to other acquisitions on the three sheeps.
To appreciate more quantitatively the benefit of UMI, the ULM density maps are superimposed on MRA, which is considered as the gold standard for observing large vascular structures in the brain. After UMI, the ULM density maps match more closely with the vessels revealed by MRA. In the coronal slice (Fig. 7A,C), UMI allows a more complete and brighter reconstruction of the vascular network (Fig. 7C) that matches in depth with the Willis polygon revealed by MRA, especially in the anterior cerebral arteries that deviate from the polygon (green and blue arrows). In the sagittal section (Fig. 7B,D), ULM provides a better adequation between MRA and ULM images of venous system (yellow arrows). Besides the repositioning of large vessels, the increase in the total SNR of the ULM volume and the appearance of small vessels provided by UMI are noteworthy compared to the initial ULM image but also with respect to MRA that only provides an image of the largest vessels.
@@ DISCUSSION
UMI approach for aberration compensation is perfectly complementary with the alternative route proposed recently by Robin et al. (?) that exploited bubbles as ultrasound guide stars for trans-cranial adaptive focusing. To some extent, UMI is more robust since it directly synthesizes virtual guide stars from speckle (?) and can work whatever the concentration of bubbles. However, in the current realization, UMI also showed some limits. First, it requires the prior recording of a high-dimension reflection matrix. Second, in the present case, it only provided the depth-dependence of aberration phase laws and was not able to capture the lateral variations of aberrations. Last but not least, it did not compensate for the multiple reverberations inside the skull that prevented us from brain imaging right below the skull.
@@
A low-frequency matrix array probe (Vermon, Tours, France), as described in Tab. 2, is placed on the shaved head of the sheep directly over the crest of the frontal bone. 3D ultrasound sequences were then recorded with a Vantage 256 4-to-1 multiplex system (Verasonics UTA 1024-MUX adapter, Kirkland, USA). The ultrasound waves were generated synthetically in this way so that the reflected wavefronts corresponding to each individual ultrasound wave emission could be reassembled in post-processing (?, ?). The successive ultrasound imaging sequences of the entire sheep experiment are summarized in Fig. 1.
@@
A UMI sequence is then performed. It consists in the acquisition of the reflection matrix using a set of 324 synthetic spherical waves whose associated virtual sources are located along a sphere radius mm: , where the vector indicates the position of the corresponding virtual sources (see Supplementary S1]) and refers to the position of the transducers. The insonification of the medium is made with diverging waves associated with virtual sources located above the probe (?) such that:
with the angle associated with each virtual source. The angular range spans from -30 to 30 with an angular pitch .
Finally, a hybrid ultrasound localization microscopy (ULM) sequence is performed. Three cylindrical waves were emitted simultaneously from the entire probe and two spherical waves were decomposed by panels to achieve a volume rate of 209 Hz (?, ?). Since only a few insonifications are used, the ULM acquisition consists of a partial reflection matrix that is acquired at several points in time so that the reflected wavefronts are stored in a second matrix . During this acquisition, microbubble injection boluses (Sonovue, Bracco, Italy) were injected, i.e. from to ml/min, during 2.5 to 6.5 minutes.
A total of 1 to 6 acquisitions were performed per sheep. The experimental conditions are summarized in Table S1. It should be noted that, for some acquisition, the UMI sequence is performed after the ULM sequence and not before, but this does not change the principle of the method.
@@
(i) Projecting onto the transducer basis () either at input or output using a homogeneous propagation model:
@@
(12) -
(ii) Isolate the distorted part of each wavefront by subtracting the geometric component that would be ideally obtained in absence of aberrations (i.e. constant speed of sound ). The result is the input and output distortion matrices which are expressed as follows:
@@
(14) with , the normalized transmission matrix such that
(iii) Exploit the angular memory effect by computing the correlation matrices of wave-front distortions,
@@
(iv) A numerical iterative phase reversal algorithm is performed to extract an estimation of aberration laws by solving the following equation:
(17) The complexity of aberration laws is quantified by the Strehl ratio that can be computed as follows (?):
@@
The whole process is repeated two times at input and output to improve gradually the focusing process. The confocal volume (Eq. 5) displayed in Fig. 4f and the map of RPSFs (Eq. 7) shown in Fig. 4d are extracted from the diagonal coefficients of this corrected focused reflection matrix.
@@
(ii) A high-pass time filter is applied to the set of images by means of a two frames sliding average window. Thereby, the static component of the ultrasound image (i.e tissues) is removed while the dynamic part of the image (moving microbubbles) is enhanced.
(iii) The microbubbles were then detected in each volume by localizing regional maxima on each image . The center of each PSF was sub-localized using the radial symmetry method (?, ?, ?), i.e. by assuming a PSF Gaussian shape and by applying a gradient algorithm. Microbubbles with a signal-to-noise ratio (SNR) below 10.5 were discarded. This SNR was calculated as the intensity at the center of the microbubbles divided by the mean intensity of the neighboring voxels over a 5-mm-wide patch. The detected bubbles and their localization are superimposed to one cross-section section of the ultrasound image in Figs. 5a,b.

### v2

# Ultrasound matrix imaging for 3D transcranial in-vivo localization microscopy

@@
Ultrasound matrix imaging for 3D transcranial in-vivo localization microscopy
Transcranial ultrasound imaging is usually limited by skull-induced attenuation and high-order aberrations. By using contrast agents such as microbubbles in combination with ultrafast imaging, not only can the signal-to-noise ratio be improved, but super-resolution images down to the micrometer scale of the brain vessels can be obtained. However, ultrasound localization microscopy (ULM) remains impacted by wave-front distortions that limit the microbubble detection rate and hamper their localization. In this work, we show how ultrasound matrix imaging (UMI), which relies on the prior recording of the reflection matrix, can provide a solution to those fundamental issues. As an experimental proof-of-concept, an in-vivo reconstruction of deep brain microvessels is performed on three anesthetized sheeps. The compensation of wave distortions is shown to drastically enhance the contrast and resolution of ULM. This experimental study thus opens up promising perspectives for a transcranial and non-ionizing observation of human cerebral microvascular pathologies, such as stroke.
@@ INTRODUCTION
To solve the fundamental problem of aberrations, adaptive focusing was introduced years ago in astronomy to compensate for phase distortions undergone by starlight when it passes through the atmosphere (?, ?, ?, ?). This idea was then transposed to ultrasound imaging by tailoring a focusing law that either optimizes an image metrics (?, ?) or maximizes the spatial correlations of back-scattered echoes (?, ?, ?, ?, ?). In the specific case of ULM, several methods have been proposed to correct the aberrations caused by the skull. A first strategy (?) is based on deep learning algorithms trained with numerical simulations. Although it gives impressive in-vivo results for small animals, it is hampered by the specificity of the training set, so its transfer to other experimental configurations (probe, organ, SNR) may require tedious training procedures. The second method (?) uses microbubbles as ultrasound guide stars but suffers several drawbacks. On the one hand, microbubbles should be detected and isolated enough, which can be extremely difficult to assess because of the strongly distorted focal spot across the skull. On the other hand, the number of detected micro-bubbles should be large enough for self-averaging in each isoplanatic patch. Last but not least, the demonstration was limited to 2D imaging using a 1D phased array probe, while skull heterogeneities are 3D distributed and require a bi-dimensional control of the ultrasonic wave-field to properly compensate for such high-order aberrations (?).
Interestingly, inspired by previous seminal works (?, ?, ?), a matrix approach of ultrasound imaging has been recently developed to overcome aberrations without any guide star (?, ?). Based on the recording of the reflection associated with the ultrasonic probe and the linearity of the wave equation, it allows the simulation of wave focusing in post-processing and an iteration of this process towards local adaptive focusing laws in ultrasonic speckle (?). An optimized contrast and close-to-ideal resolution is then obtained for each voxel of the ultrasound image. Ultrasound Matrix Imaging (UMI) has been successfully applied to transcranial 3D imaging of a head phantom (?).
The goal of this paper is to demonstrate the interest of UMI for in-vivo 3D brain imaging. In particular, we will show how: (i) UMI can be used to quantify aberrations and multiple scattering in in-vivo transcranial imaging; (ii) UMI can exploit static speckle from brain tissue to tailor complex 2D focusing laws in a strong aberration and multiple scattering regime; (iii) ULM can leverage these focusing laws despite motion to increase the detection rate of microbubbles and improve their localization; (iv) UMI can be fruitfully combined to ULM in order to provide a three-dimensional, artifacts-free, super-resolved and contrasted image of brain vessels.
@@ RESULTS
In this paper, we demonstrate transcranial brain imaging by overcoming the major challenges of: (i) skull bone aberration with 3D UMI; (ii) diffraction-limited resolution with the super-resolution feature provided by ULM. As a proof-of-concept, the same experimental protocol (Fig. 1) has been applied to three anesthetized sheeps (Methods). A array of transducers is placed on the shaved head of each sheep directly over the crest of the frontal bone (Fig. 1B). The probe is driven in 1-2 frequency bandwidth in order to limit the attenuation by skull bone (?). The first ultrasound sequence consists in recording a high-dimension reflection matrix (Fig. 1C) that contains the scattered wave-field recorded by each transducer of the probe as a function of the echo time for a set of 324 incident diverging waves. This illumination basis is associated with a set of virtual transducers placed behind the probe (see Methods and Fig. S1). It maximizes the signal-to-noise ratio across the targeted field-of-view (?). The number of incident waves is dictated by our choice of the maximal angular range spanning from to (Methods).
@@
where and are the coordinates of the input and output focal points, respectively. and are the expected travel times for the wave to go from the source/detector to the input/output focusing point (Methods, Eq. 4). The result of Eq. 1 is a focused reflection matrix whose coefficients probe the cross-talk between a virtual source and detector at lateral positions and , respectively, both located at the same depth inside the medium (Fig. 2A). This matrix enables a local quantification of the focusing quality and multiple scattering by providing a local point spread function in reflection (RPSF) (?, ?, ?) (Methods, Eq. 7). Figure 3B display maps of RPSFs at depth mm for each sheep. In the single scattering regime and in absence of aberrations, the RPSF would exhibit a confocal diffraction-limited peak. On the contrary, the various RPSFs here display a distorted focal spot induced by wave velocity heterogeneities on top of an incoherent background due to multiple scattering events induced by the skull. Strikingly, whatever the imaging depth, the RPSF looks more distorted and less contrasted for sheep n6 than for sheep n4 and 5 (Fig. 3B).
On the one hand, the aberration level can be quantified by investigating the spatial extent of the RPSF (?, ?) (Fig. S4). The depth evolution of is shown in Fig. 3C. A general trend is a degradation of the resolution with depth as predicted by diffraction theory but also a more important aberration level in sheep n6. This difference of behavior can be understood by looking at the micro-CT of each skull (Fig. 3A). While the skulls of sheep n4 and 5 show a regular thickness ( mm and mm, respectively), skull n6 displays a more irregular shape: mm (?). This difference of morphology between the three skulls could explain the higher degree of aberrations observed in sheep n6.
On the other hand, the multiple scattering rate can be estimated from the incoherent background of the RPSF (Methods, Eq. 9). It appears that the multiple scattering rate is far from being negligible even after the beamforming process since it reaches a value of 30 % in the mm-region for sheep n6 (Fig. 3D). As the aberration level, the multiple scattering rate is larger for this sheep over almost the whole depth range. This feature can be understood by looking at the internal structure of the bone. Indeed, the diploë volume ratio is much higher in sheep n6 (%) than in sheeps n4 and 5 (% and %, respectively) (?). This high degree of heterogeneity seems to account for the higher multiple scattering rate observed in sheep n6.
@@
The complex speed-of-sound distribution highlighted by Fig. 3B hampers ultrasound brain imaging. The confocal image of the brain, that can be extracted from the diagonal coefficients of (Methods, Eq. 5), is blurred by the skull heterogeneities (Fig. 4A). Fortunately, an optimized contrast and diffraction-limited resolution can be recovered using the distortion matrix concept (?, ?, ?). This matrix contains the wave distortions exhibited by each reflected wave-front in the transducer basis (Fig. 2C). The choice of this basis is dictated by the fact that most aberrations are induced by the skull heterogeneities, i.e in the vicinity of the probe. An iterative phase reversal (IPR) process is then applied to the distortion matrix to provide an estimation of an aberration phase law, , at each depth (Methods, Fig. 2D and Movie S1). Examples of aberration phase laws are displayed in Fig. 4C. They show a complex feature, as quantified by the Strehl ratio that provides the ratio between the energy at focus with and without aberrations (?) (Methods, Eq. 18). For instance, the aberration law measured at depth mm exhibits a Strehl ratio . This extremely low value illustrates the detrimental impact of skull heterogeneities on the focusing process. A more detailed quantification of the aberration magnitude is provided in Section S3 by evaluating the modulus of the aberration transfer function. As expected from the previous RPSF analysis (Fig. 3), the aberration magnitude is found to be 30 % larger in sheep n6 compared to sheep n4 and 5.
The aberration phase laws can be leveraged to compensate for the wave distortions exhibited by the focused reflection matrix (Methods). The resulting confocal image is displayed in Fig. 4E. It exhibits a clear improvement compared to the initial confocal image (Fig. 4A), with a contrast gain of 8 dB for the deepest bright scatterers ( and mm). The efficiency of the correction process can be assessed by comparing the maps of RPSFs before (Fig. 4B) and after (Fig. 4D) aberration compensation. The spatial resolution is improved by a factor between 2 and 3 and the multiple scattering background is decreased by more than 10 dB beyond mm (Fig. S5). Nevertheless, the obtained RPSFs remain imperfect in Fig. 4D. High-order aberrations subsist because of their anisoplanicity. Last but not least, even though the confocal image shows a clear improvement, the diffraction-limited resolution remains limited by the probe size : .
@@
A deconvolution of the dynamic image with a theoretical PSF and the application of a detection threshold provides a localization of bubbles in the field-of-view (Methods). The position of the detected bubbles are highlighted with red points in Fig. 5A and B. While several microbubbles are wrongly assigned to the high side-lobes of each distorted focal spot in Fig. 5A, UMI drastically reduces this number of false detection in Fig. 5B by restoring a satisfying PSF, in particular in the colored squares. Despite this reduction of false alarms, the number of localized microbubbles is increased. Figure 5C illustrates this assertion by showing, for sheep n, the higher number of detected bubbles over time with UMI. The different peaks corresponds to each injection of microbubbles.
To further reduce the false alarm rate and better appreciate the effect of UMI, microbubbles can be tracked over time (Methods). In absence of aberration correction, the focal spot is distorted and bubbles cannot be detected on each frame during their journey across the field-of-view. The bubble trajectory is then subdivided into shorter tracks. As a consequence, the track length is a relevant observable to demonstrate the benefit of the aberration correction process (?). Fig. 5D shows the increase of detected tracks provided by UMI as a function of their length. UMI significantly increases the number of detected tracks by a factor ranging from 40 to 250% according to the track length. Table 1 shows the same observable for different minimum path lengths. The general trend is the same for each acquisition, thereby showing the robustness of UMI and its benefit in each imaging configuration. This result is crucial for ULM images as the longest tracks are the most reliable to delineate small vessels.
@@
| 2 frames | >3 frames | >5 frames | >10 frames | >30 frames | ||
@@
However, the skull heterogeneities also induce artifacts such as the duplication of one artery that can be noticed at the bottom of the conventional ULM image (Figs. 6A and C). This type of artifact has already been observed in transcranial imaging in both mouse (?) and human (?). Strikingly, the focusing laws extracted by UMI (Fig. 4C) allow the removal of such artifacts as highlighted by the corrected ULM images in Figs. 6B and D. More generally, the higher detection rate and correct repositioning of micro-bubbles lead to a better contrast for ULM as highlighted by the vessel network on top of the Willis circle (25-35 mm) that appears much brighter (+6 dB) and better resolved. A similar resolution and contrast improvement can be observed for the other 3D ULM images shown in Fig. S9 and corresponding to other acquisitions on the three sheeps.
To appreciate more quantitatively the benefit of UMI, the ULM density maps are superimposed on MRA, which is considered as the gold standard for observing large vascular structures in the brain. After UMI, the ULM density maps match more closely with the vessels revealed by MRA. In the coronal slice (Fig. 7A,C), UMI allows a more complete and brighter reconstruction of the vascular network (Fig. 7C) that matches in depth with the Willis polygon revealed by MRA, especially in the anterior cerebral arteries that deviate from the polygon (green and blue arrows). A zoom on the right anterior cerebral artery is displayed in Fig. S8 and highlights the drastic gain in contrast provided by ULM compared to MRA around the Willis circle. This benefit of UMI can be quantified by a loss of image entropy (?) increased by a factor 5 with UMI.
In the sagittal section (Fig. 7B,D), ULM provides a better adequation between MRA and ULM images of venous system (yellow arrows). This observation can be quantified by a computing structural similarity index between MRA and ULM images. This similarity metrics is increased by a factor 2 with UMI (Fig. S7). Besides the repositioning of large vessels, the increase in the total SNR of the ULM volume and the appearance of small vessels provided by UMI are noteworthy compared to the initial ULM image but also with respect to MRA that only provides an image of the largest vessels (Fig. S8).
Finally, note that the volume rate (209 Hz) of our system limits the ability of the ULM pipeline to track microbubbles traveling in the larger cerebral vessels and to provide an appropriate flow velocity map (?). In the future, a 3D ultrasound scanner with a higher volume rate could alleviate this issue and provide additional physiological information.
@@ DISCUSSION
UMI approach for aberration compensation is perfectly complementary with the alternative route proposed recently by Robin et al. (?) that exploited bubbles as ultrasound guide stars for trans-cranial adaptive focusing. To some extent, UMI is more robust since it directly synthesizes virtual guide stars from speckle (?) and can work whatever the concentration of bubbles as well as in stronger aberration and multiple scattering conditions. Typically, we were not able to apply the approach proposed in Ref. (?) due to the low quality of the ULM data. Indeed, the latter method requires the presence of bubbles whose associated focal spot displays a correlation degree larger than 0.6 with the ideal diffraction-limited PSF. Here, the normalized Strehl ratio, whose value is a maximal bound for , ranges from 0.1 to 0.4 (Section S3). Such an aberration level makes difficult the use of microbubbles as guide stars in an adaptive focusing scheme. In Section S6, we also show how UMI provides a flexible benchmarking framework to show the benefits of the iterative phase reversal process (?) compared with conventional aberration correction methods in speckle such as iterative time reversal (?) or SVD beamforming (?, ?).
Nevertheless, in the current realization, UMI also showed some limits. First, it requires the prior recording of a high-dimension reflection matrix. Second, in the present case, it only provided the depth-dependence of aberration phase laws and was not able to capture the lateral variations of aberrations. Last but not least, it did not compensate for the multiple reverberations inside the skull that prevented us from brain imaging right below the skull. Those limits are explained by the fact that high-order aberrations and reverberations are associated with extremely small isoplanatic patches. This feature limits the efficiency of the spatial averaging process required for the synthesis of a coherent guide star in matrix imaging (Fig. 2).
@@
A low-frequency matrix array probe (Vermon, Tours, France), as described in Table 2, is placed on the shaved head of the sheep directly over the crest of the frontal bone. 3D ultrasound sequences were then recorded with a Vantage 256 4-to-1 multiplex system (Verasonics UTA 1024-MUX adapter, Kirkland, USA). The 1024 elements were connected to the 256 channels of the ultrasound scanner using a 4-to-1 multiplexer (Verasonics UTA 1024-MUX Adapter). The probe is therefore divided into four panels, the multiplexer only allowing emission with one panel and receiving with another (?, ?). For the reflection matrix acquisition, a set of 16 emissions had to be performed for any incident wave-field in order to record the responses between each panel at emission and reception. Those recordings are then reassembled in post-processing to get the full reflected-wave-front associated with each synthetic incident wave-field. For the ULM sequence, this constraint has been partially mitigated since it is possible to emit with the four connected panels at the same time, provided that the emission waveform and delays were identical along the elevation axis (?, ?). As detailed further, the use of incident cylindrical wave-fronts was therefore used to reach a sufficient frame rate for ULM. The successive ultrasound imaging sequences of the entire sheep experiment are summarized in Fig. 1.
@@
A UMI sequence is then performed. It consists in the acquisition of the reflection matrix using a set of 324 synthetic spherical waves whose associated virtual sources are located along a sphere radius mm: , where the vector indicates the position of the corresponding virtual sources (see Supplementary S1) and refers to the position of the transducers. The insonification of the medium is made with diverging waves associated with virtual sources located above the probe (?) such that:
with the angle associated with each virtual source. The angular range spans from to , with . The angular pitch is governed by the probe extension: , with , the central wave length. This angular range and angular pitch dictate the number of incident diverging waves: .
Finally, a hybrid ultrasound localization microscopy (ULM) sequence is performed. To limit the number of incident waves, only the azimuth and elevation shifts of (0,0); (-5, 0); (5, 0); (0, -5); (0, 5) degrees are considered. As stated above, the 4-to-1 multiplexer imposes a sequence of 16 insonifications to record the response of the medium for each virtual source. Such a sequence is prohibitive for high-frame rate imaging. However, this condition can be relaxed by only recording the responses between neighboring panels, thereby restricting the number of insonifications to 10 for each virtual source (?). Moreover, the four panels can be connected together for the emission if the delays are similar in the elevation direction. For such an incident cylindrical wave-front, only 4 successive transmissions are required for the 4 reception apertures. However, full cylindrical waves could only be induced when the angle is 0 in the azimuthal direction. In the direction, waves still needed to be split in multiple emission-reception. Consequently, an hybrid sequence was chosen to optimize the frame rate while maintaining a sufficient angular diversity for the acquisition:
Three cylindrical waves were emitted sequentially from the entire probe and had to be repeated four times to record the reflected wave-front on each panel.
Two spherical waves (virtual source) were decomposed by panels, corresponding to 10 emissions.
The full ULM sequence thus corresponds to 32 emissions to achieve a volume rate of 209 Hz (?, ?). Since only a few insonifications are used, the ULM acquisition consists of a partial reflection matrix that is acquired at multiple times so that the reflected wavefronts are stored in a second matrix . The sparse feature of the recorded reflection matrices implies an ultrasound image of lower quality compared to a full measurement of the -matrix. Both the contrast and, to a lessen extent, the resolution of the beamformed image are degraded. However, those detrimental effects are compensated, at least partially, by the echogenicity of the bubbles for the contrast and by the localization feature that allows the resolution to be no longer limited by diffraction but by the signal-to-noise ratio (?). During this acquisition, microbubble injection boluses (Sonovue, Bracco, Italy) were injected, i.e. from to ml/min, during 2.5 to 6.5 minutes.
A total of 1 to 6 acquisitions were performed per sheep. The experimental conditions are summarized in Table S1. The raw ULM data used in the paper corresponds to the ULM data presented in Ref. (?). It should also be noted that, for some acquisition, the UMI sequence is performed after the ULM sequence and not before, but this does not change the principle of the method.
@@
(i) We first project onto the transducer basis () either at input or output using a homogeneous propagation model:
@@
(12) As sketched in Fig. 2B, this dual reflection matrix contains the wave-fronts associated with virtual sources corresponding to the focusing points . Those wave-fronts correspond to the sum of a geometric parabolic component related to to the position of each focusing point and of a distorted component resulting from the mismatch between the propagation model and the real speed-of-sound distribution. Interestingly, this distorted component displays strong correlations because each reflected wave-front goes through the same part of the skull, the probe being placed at the skull surface. This is the so-called angular (or tilt-tilt) memory effect (?) that we will exploit in the following.
(ii) To do so, we isolate the distorted part of each wavefront by subtracting the geometric component that would be ideally obtained in absence of aberrations (i.e. constant speed of sound ). The result is the input and output distortion matrices which are expressed as follows:
@@
(14) where the symbol stands for the Hadamard (element wise) product and with, , the normalized transmission matrix such that
In the focused basis, the operation described in Eqs. 13 and 14 amounts to a de-scan of each virtual source at the origin of each focal plane, as sketched in Fig. 2C. In the transducer basis, it consists in a realignment of associated wave-fronts, thereby highlighting the angular correlations between wave distortions.
(iii) We then exploit this angular memory effect by computing the correlation matrices of wave-front distortions,
@@
(iv) A numerical iterative phase reversal (IPR) algorithm (?) is performed to extract an estimation of aberration laws by solving the following equation:
(17) The IPR process, sketched in Fig. 2D, allows the synthesis of a coherent guide star from the set of virtual sources . The associated wave-front is the resulting aberration phase law . The complexity of each aberration law can be quantified by the Strehl ratio which is defined as follows (?) (see also Section S3):
@@
The whole process is repeated two times at input and output to improve gradually the focusing process. The whole UMI process is summarized by the flow chart displayed in Fig. S2. The confocal volume (Eq. 5) displayed in Fig. 4f and the map of RPSFs (Eq. 7) shown in Fig. 4d are extracted from the diagonal coefficients of this corrected focused reflection matrix.
@@
(ii) A high-pass time filter is applied to the set of images by means of a two frames sliding average window:
Albeit crude, this simple operation allows the removal of most of the slowly-varying component of the ultrasound image (i.e tissues) while enhancing the dynamic part of the image (moving microbubbles).
(iii) The microbubbles were then detected in each volume by localizing regional maxima on each image . The center of each PSF was sub-localized using the radial symmetry method (?, ?, ?), i.e. by assuming a PSF Gaussian shape and by applying a gradient algorithm. Microbubbles with a signal-to-noise ratio (SNR) below 10.5 dB were discarded. This SNR was calculated as the intensity at the center of the microbubbles divided by the mean intensity of the neighboring voxels over a 5-mm-wide patch. The detected bubbles and their localization are superimposed to one cross-section section of the ultrasound image in Figs. 5a,b.